OpenStudy (anonymous):

What are the approximate solutions of 2x2 − 4x = 1 rounded to the nearest hundredth? help please

OpenStudy (anonymous):

@sara17

OpenStudy (anonymous):

i think first we should solve equation and find x

OpenStudy (anonymous):

ok sounds good to me

OpenStudy (anonymous):

do u know how to solve it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so can you explain

OpenStudy (anonymous):

ofcource.. first move all variables and numbers to one side of the equation: 2x^2-4x-1=0 then try to remove x's coefficient so divid eaquation by2: x^2-2x-1/2=0 then try to make this equation square to solve it easily so add andsubtract 1 to the equation so: x^2-2x+1-1-1/2=0 (x^2-2x+1)-3/2=0 (x-1)^2=3/2 so radical: x-1=+-(3/2) x=5/2 x=-1/2

OpenStudy (anonymous):

thats not one of the choices

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

ok look x ≈ −0.23 and x ≈ 2.23 x ≈ 1.71 and x ≈ 0.29 x ≈ −0.45 and x ≈ 4.45 No real solutions

OpenStudy (anonymous):

oh yes.. these are approximations..

OpenStudy (anonymous):

\[2 x^2-4 x-1 \]cannot be factored. Refer to the attached Mathematica calculation.

OpenStudy (anonymous):

ok? im confused

OpenStudy (anonymous):

sorry i made mistake sqrt(3/2)=\[\sqrt{3/2}\] so x=\[\sqrt{3/2}\]+1 x=-\[\sqrt{3/2}\]+1 then what's robtobey said was true.. sorry for that mistake

OpenStudy (anonymous):

its ok thanks guys